Prove that ((x+y+z)/3)^(x+y+z) ≤ x^x y^y z^z ≤ ((x^2+y^2+z^2)/(x+y+z))^(x+y+z)
1
Expert's answer
2020-05-25T15:58:36-0400

Due to the domain of the function "a^x," namely "a > 0", we obtain that "x,y,z>0".

Let's prove the first inequality:

"(\frac{x+y+z}{3})^{(x+y+z)}\le x^xy^yz^z". By dividing both parts by "A^{x+y+z} = (\frac{x+y+z}{3})^{(x+y+z)}", we obtain the following:

"(\frac{x}{A})^{x}(\frac{y}{A})^{y}(\frac{z}{A})^{z} \ge 1." Since "A>0", we can write the following:

"(\frac{x}{A})^{\frac{x}{A}}(\frac{y}{A})^{\frac{y}{A}}(\frac{z}{A})^{\frac{z}{A}} \ge 1." Let's consider the following function: "u(t,u,v) = t^tu^uv^v", but where "t+u+v = 3" , respectively to our case. Let's find its local minima and maxima of this function(to prove it is "\ge 1, t,u,v \in R" ). Using the method of Lagrange multipliers we obtain the following system:

"\begin{cases}
t^tu^uv^v(\ln{t} + 1) - \lambda = 0 \\
t^tu^uv^v(\ln{u} + 1) - \lambda = 0 \\
t^tu^uv^v(\ln{v} +1) - \lambda = 0 \\
t+u+v=3
\end{cases}" . It is obviously inferred from the first 3 equations, that "t = u =v". Using the 4th equation, we obtain that "t = u = v = 1" . Let's figure out if this is a local maximum or minimum. Using the Hessian matrix method we obtain that it is local minimum. But, the inequality is true at "(1,1,1)". Since the function had the minimum at "(1,1,1)", we obtain that initial inequality is true for every "(t, u, v)", and, hence, for each "(x,y,z) \in R^3".


Let's now prove the 2nd inequality:

"x^xy^yz^z \le (\frac{(x^2 + y^2 + z^2)}{x+y+z})^{x+y+z}" . Using the Weighted AM-GM Inequality, for "\omega_1 = x, \omega_2 = y, \omega_3 = z," we obtain: "(x^xy^yz^z)^{\frac{1}{x+y+z}} \le \frac{x\cdot x + y\cdot y + z \cdot z}{x+y+z} = \frac{x^2 + y^2 + z^2}{x+y+z}" . Since "x+y+z > 0" , raising both parts of the inequality to the "x+y+z" power, we obtain the following:

"x^xy^yz^z \le (\frac{x^2 + y^2 +z^2}{x+y+z})^{x+y+z}" .